• Find the area of ​​a triangle constructed on vectors а‾ = ı‾ + ȷ‾ , ̅b‾=2ȷ‾.
1
Expert's answer
2022-06-16T10:33:46-0400
"\vec{a}\times\vec{b}=\begin{vmatrix}
 \vec{i} & \vec{j}& \vec{k} \\
 1 & 1 & 0 \\
 0 & 2 & 0 \\
\end{vmatrix}=\vec{k}(2-0)=2\vec{k}""Area=A=\dfrac{1}{2}|\vec{a}\times\vec{b}|=\dfrac{1}{2}(2)=1({units}^2)"

1 square unit.